OpenStudy (anonymous):

solve 2 minus the squareroot of 2 minus x = x + 6

OpenStudy (anonymous):

\[2 - \sqrt{2} - x = x+6\] \[2x = -4 + \sqrt{2}\] \[x = -2 +\sqrt{2}/2 = -1.292\]

myininaya (myininaya):

\[2-\sqrt{2}-x=x+6\] \[2x=2-6-\sqrt{2}=-4-\sqrt{2}\] \[x=\frac{-4-\sqrt{2}}{2}=\frac{-4}{2}-\frac{\sqrt{2}}{2}=-2-\frac{\sqrt{2}}{2}\]

OpenStudy (anonymous):

^ I really need to think about retirement :D

myininaya (myininaya):

shhh daniel your awesome